OpenStudy (anonymous):

Somebody PLEASE help me!! i will medal! Roberto wants to estimate the answer to this problem to see if his answer is reasonable. 5849 – 2621 Which expression shows these numbers correctly rounded to the nearest hundred? A. 5900 – 2600 B. 5800 – 2600 C. 5900 – 2700 D. 5800 – 2700

OpenStudy (anonymous):

@YoungStudier

OpenStudy (youngstudier):

What do you think?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

sorry im bad at rounding

OpenStudy (youngstudier):

B is correct. :) Good job!

OpenStudy (anonymous):

thanks!

OpenStudy (youngstudier):

No problem, have any more?

OpenStudy (anonymous):

yeah one more

OpenStudy (youngstudier):

Post it here, then. :)

OpenStudy (anonymous):

Julia solved this problem. Peter had a strip of leather 58.8 cm long. He cut the strip of leather into 6 equal-length pieces. How long was each piece of leather? Julia said the answer was 0.98 cm. Is Julia's answer reasonable? A. No, because 60 • 6 = 360. B. No, because 60 ÷ 6 = 10. C. No, because 60 + 6 = 66. D. Yes

OpenStudy (anonymous):

i think this one is also B

OpenStudy (anonymous):

but im not sure

OpenStudy (youngstudier):

I'm actually not entirely sure for this one. O_O 58.8/6 = 9.8 I don't know how they would get 0.98........

OpenStudy (youngstudier):

@JackofallTradez

OpenStudy (anonymous):

@Missiey

OpenStudy (youngstudier):

@PrincessHush @CallMeKiki

OpenStudy (youngstudier):

@TheEdwardsFamily

OpenStudy (youngstudier):

*Sigh* I would just go with what you think best. I'm not sure for this one, sorry.

OpenStudy (anonymous):

its okay, thanks anyways!

OpenStudy (youngstudier):

You're welcome.